# A Non-Relativistic Limit for Heterotic Supergravity and its Gauge Lagrangian

### v3

@@ Abstract
We show that the $D=10$ heterotic supergravity in a non-relativistic (NR) limit has a finite Lagrangian due to non-trivial cancellations of divergent parts arising from the Chern-Simons terms in the curvature of the $\hat{B}$-field and the Yang-Mills Lagrangian. This is similar to what happens in bosonic supergravity between the Ricci scalar, $\hat{R}$, and the $-\frac{1}{12}\hat{H}^{2}$ term, under the same limit. In this work we present the explicit form of the gauge transformations and curvatures after considering a NR limit and we compute the full finite Lagrangian in its covariant form. Particularly, we address both the ungauged part and the gauged part, and we write it using covariant derivatives and curvatures for the gauge fields. These results constitutes the first computation of a finite Lagrangian for heterotic supergravity under these kinds of limits. As an interesting feature of the expansion implemented, the Green-Schwarz mechanism for the two-form can be trivialized and terms equivalent to Chern-Simons contributions naturally arise from the previous property.
@@ 1 Introduction
In the last decade there has been huge progress made in the study of the different formulations of non-relativistic string theory [1]-[3] and its supergravity limit [4]-[17]  [^1] See [18]-[20] for reviews on this topic.. The low energy limit of bosonic string theory is given by the universal NS-NS supergravity, whose fundamental fields $\hat{g}$, $\hat{B}$, $\hat{\phi}$ admit a non-relativistic expansion of the form [13],
@@
where $\mu,\nu=1,\dots,10$ are space-time indices and we have split the flat index $\hat{a}=(a,a^{\prime})$ where $a,b,...=0,1$ are the transverse directions and $a^{\prime},b^{\prime},...=2,\dots,9$. The quantities in the previous expansion obey the following Newton-Cartan relations,
@@
While, apriori, this expansion produces divergent tensorial quantities, i.e., Riemann tensor and its traces, H strength, etc., the NS-NS supergravity Lagrangian,
@@
remains finite due to a curious cancellation between the divergent part of the Ricci scalar and the divergent part of the $\hat{H}^{2}$ term [13].
In the case of the bosonic sector of heterotic supergravity, we expect a similar cancellation when including the gauge field $\hat{A}_{\mu}^{i}$, $i=1,\dots,N$ and $N=496$ which is the dimension of the heterotic gauge group. Given a particular expansion for the gauge field, the inclusion of Chern-Simons terms in the curvature of the $\hat{B}$-field should produce divergences that must be compensated by the Yang-Mills Lagrangian. While there are several candidates for the expansion of the A-field in the heterotic scenario [21]-[22], there is still no explicit computation of the finite form of the heterotic Lagrangian under this limit. The main reason of an explicit computation of this Lagrangian is that exact results under the NR limit usually contain large contributions when they are not written in covariant form. We expect that the NR limit of the heterotic supergravity admits a rewriting in terms of $O(D,D+N)$ multiplets. T-duality is an exact symmetry of string theory and the low energy limit of the theory should be able to written in terms of a generalized metric and dilaton. In other words, we expect that the formalism of Double Field Theory [23]-[26] could be able to rewrite the theory, when the fundamental degrees of freedom are parametrized using a non-Riemannian ansatz [27]-[30].
On the other hand, we expect that the vielbein remains gauge invariant while the $\hat{B}$-field inherits a non-covariant gauge transformation (Green-Schwarz mechanism), similarly to what happens in the relativistic case, i.e.,
@@
$\delta_{\hat{\lambda}}\hat{g}_{\mu\nu}$ $=$ $0\,,$ (1.11)
where $\hat{\lambda}$ is a gauge parameter. Due to the previous restrictions, we will expand the gauge field following the proposal given in [22]
@@
where $v^{\pm}=\frac{1}{\sqrt{2}}(v^{0}\pm v^{1})$. The expansion (1.12) was constructed in such a way that the fundamental degrees of freedom of DFT remains convergent after taking the NR limit in the double geometry. In this way, we are sure that the NR heterotic supergravity can be rewritten in terms of $O(D,D+N)$ multiplets.
@@
remains finite. We will also present the explicit form of the gauge transformations and the heteorotic supergravity Lagrangian in its covariant form.
@@ 1.1 Organization of the paper
This paper is organized as follows: In the next section we review the cancellation of divergences in bosonic supergravity and we show the cancellation of divergencies in heterotic supergravity. In section 3 we study the gauge transformation rules of the fundamental fields of the theory upon considering the NR limit and we construct a suitable covariant derivative and curvatures for all the fields that transform in a non-covariant fashion. In section 4 we compute the finite Lagrangian of the theory. We will write not only the gauge part using new curvature for the gauge fields, but also the ungauged part of the Lagrangian [13] using new covariant structures which explicit show the invariance under infinitesimal diffeomorphisms. Finally, in section 5, we present a discussion and an outlook.
@@ 2 Cancellation of divergences
When one expands the action (1.7), both the $\hat{R}$ and $-\frac{1}{12}\hat{H}^{2}$ terms produce $c^{2}$ contributions. The same happens when we analyze (1.13), but this time the gauge compensation of $-\frac{1}{12}\bar{H}^{2}$ is given also by the $-\frac{1}{4}\hat{F}^{2}$ term. We will analyze both scenarios separately.
@@ 2.1 Ungauged case (review)
Let’s review the cancellation of divergences in the bosonic supergravity Lagrangian. To do so, we define
@@
$\hat{R}^{(2)}=-\frac{c^{2}}{2}\eta_{ab}\partial_{\mu}{\tau_{\nu}{}^{a}}\partial_{\rho}{\tau_{\sigma}{}^{b}}(h^{\mu\sigma}h^{\nu\rho}-h^{\mu\rho}h^{\nu\sigma})\,.$ (2.2)
So far we have just considered the expansion of the metric tensor (1.1) and its inverse (1.2). Therefore, for backgrounds where $h^{\mu\nu}$ is constant, the divergence in the Ricci scalar becomes trivial and it is possible to define a non-relativistic supergravity without including a non-relativistic expansion in the $\hat{B}$-field. For a more general scenario we impose (1.3) and then the divergent part of $-\frac{1}{12}\hat{H}^{2}$ is given by
$-\frac{1}{12}(\hat{H}_{\mu\nu\rho}\hat{H}^{\mu\nu\rho})^{(2)}=-\frac{c^{2}}{2}\epsilon_{ab}\epsilon_{cd}\eta^{ac}\partial_{\mu}\tau_{\nu}{}^{b}\partial_{\rho}\tau_{\sigma}{}^{d}(h^{\mu\rho}h^{\nu\sigma}-h^{\mu\sigma}h^{\nu\rho})\,.$ (2.3)
Using the relation $\epsilon_{ab}\epsilon_{cd}\eta^{ac}=\eta_{bd}$ one can cancel the divergences coming from the Ricci scalar.
@@ 2.2 Gauged case
upon considering the non-relativistic expansions (1.12) and (1.2) contains a divergence of the form (see appendix 5 for intermediate steps)
@@
These terms indicate that the $A$-expansion in (1.12) does not lead to a finite Yang-Mills theory. However, the inclusion of Chern-Simons terms in the heterotic Lagrangian (1.13) compensates the divergence (2.5). Indeed, one can compute (we use CADABRA [31] for performing this test. The details are given in the appendix),
@@
At this point, we have proved that the Heterotic supergravity Lagrangian is finite, but trying to write it in non-covariant form leads to huge number of terms. For this reason, our strategy will be to write it in covariant form, making use of the diffeomorphism and gauge invariance of the action.
In the next section, we will study the form of the gauge transformation rules and we will construct a suitable covariant derivative. After that we will construct new curvatures for all the fundamental fields transforming in a non-covariant way under gauge transformations.
@@ 3 Transformations and curvatures
In this section we will focus on two symmetries: infinitesimal diffeomorphism and the non-Abelian gauge transformations associated to the $A$-field.
@@ 3.1 Covariant derivative
$=$ $\partial_{\mu}\tau_{\nu}{}^{a}+\partial_{(\mu}{\tau_{\nu)}{}^{a}}-\frac{1}{c^{2}}\partial_{(\mu}{\tau_{\nu)}{}^{a}}+T_{\mu\nu}^{a}\,,$
where $T_{\mu\nu}{}^{a}$ is given by the following covariant quantity,
$T_{\mu\nu}{}^{a}$ $=$ $\eta_{bc}\eta^{ad}\nabla_{(\mu|}{\tau_{\rho}{}^{b}}\tau_{|\nu)}{}^{c}\tau^{\rho}{}_{d}-\eta_{bc}\eta^{ad}\nabla_{\rho}{\tau_{(\mu|}{}^{b}}\tau_{|\nu)}{}^{c}\tau^{\rho}{}_{d}$ (3.2)
$-\frac{1}{c^{2}}\eta_{bc}\eta^{ad}\nabla_{(\mu}{\tau_{\rho}{}^{b}}\tau_{\nu)}{}^{c}\tau^{\rho}{}_{d}+\frac{1}{c^{2}}\eta_{bc}\eta^{ad}\nabla_{\rho}{\tau_{(\mu}{}^{b}}\tau_{\nu)}{}^{c}\tau^{\rho}{}_{d}\,.$
An interesting observation of $\nabla_{\mu}\tau_{\nu}{}^{a}$ is that this quantity does not diverge, but contains $\frac{1}{c^{2}}$ that we need to track in order to correctly write the finite Lagrangian. This means that the finite Lagrangian will evolve terms with powers of $c$, but this is just the price to pay for writing it in covariant form. The antisymmetric part of $\partial_{\mu}\tau_{\nu}^{a}$ satisfies
$\nabla_{[\mu}\tau_{\nu]}^{a}$ $=$ $\partial_{[\mu}\tau_{\nu]}{}^{a}$ (3.3)
while the symmetric part can be written as
$\partial_{(\mu}\tau_{\nu)}{}^{a}=\frac{c^{2}}{2c^{2}-1}\Big{(}\nabla_{(\mu}\tau_{\nu)}{}^{a}-T_{\mu\nu}{}^{a}\Big{)}=Q_{\mu\nu}{}^{a}\,.$ (3.4)
On the other hand the covariant derivative of $\tau^{\nu}{}_{a}$ is given by
$\nabla_{\mu}\tau^{\nu}{}_{a}=\partial_{\mu}\tau^{\nu}{}_{a}+\Gamma_{\mu\rho}^{\nu}\tau^{\rho}{}_{a}=\frac{1}{2}\partial_{\mu}\tau^{\nu}{}_{a}+S_{\mu}{}^{\nu}{}_{a}\,,$ (3.5)
where $S_{\mu}{}^{\nu}{}_{a}$ is a covariant quantity given by
$S_{\mu}{}^{\nu}{}_{a}$ $=$ $c^{2}\Big{(}\nabla_{[\rho}{\tau_{\sigma]}^{b}}\eta_{bc}\tau_{\mu}^{c}\tau^{\rho}_{a}h^{\nu\sigma}+\nabla_{[\mu}{\tau_{\rho]}^{b}}\eta_{ab}h^{\nu\rho}\Big{)}+\frac{1}{2}Q_{\mu\rho}{}^{b}(\tau^{\nu}_{b}\tau^{\rho}_{a}+\eta_{ab}h^{\nu\rho})$ (3.6)
$+\frac{1}{2}\nabla_{[\mu}{\tau_{\rho]}^{b}}(2\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d}-\tau^{\nu}_{b}\tau^{\rho}_{a}-\eta_{ab}h^{\nu\rho})+\nabla_{[\rho}{\tau_{\sigma]}^{b}}(\eta_{bc}\tau_{\mu}^{c}\tau^{\sigma}_{a}h^{\nu\rho}+\eta_{bc}\eta^{de}\tau_{\mu}^{c}\tau^{\nu}_{d}\tau^{\rho}_{a}\tau^{\sigma}_{e})$
$+\frac{1}{4c^{2}}\Big{(}Q_{\mu\rho}{}^{b}(-2\tau^{\nu}_{b}\tau^{\rho}_{a}+2\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d})+\nabla_{\mu}{\tau_{\rho}^{b}}(\tau^{\nu}_{b}\tau^{\rho}_{a}-\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d})$
$-\nabla_{\rho}{\tau_{\mu}^{b}}(\tau^{\nu}_{b}\tau^{\rho}_{a}-\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d})+4\nabla_{[\rho}{\tau_{\sigma]}^{b}}\eta_{bc}\eta^{de}\tau_{\mu}^{c}\tau^{\nu}_{d}\tau^{\sigma}_{a}\tau^{\rho}_{e}\Big{)}\,.$
Considering the Newton-Cartan relation, it is easy to prove that the ungauged Lagrangian can be written in terms of $\nabla_{\mu}\tau_{\nu}{}^{a}$, $\nabla_{\mu}\tau^{\nu}{}_{a}$ and the tensors $Q_{\mu\nu}{}^{a}$ and $S_{\mu}{}^{\nu}{}_{a}$, which also depends on the previous covariant derivatives. In other words, we do not need to compute the covariant derivative of $h_{\mu\nu}$ and $h^{\mu\nu}$ to write, for example, the Riemann tensor, the Ricci tensor or the Ricci scalar in covariant form.

### v5

@@ Abstract
We show that the $D=10$ heterotic supergravity under a non-relativistic (NR) limit has a finite Lagrangian due to non-trivial cancellations of divergent parts arising from the Chern-Simons terms in the curvature of the $\hat{B}-$field and the Yang-Mills Lagrangian. This is similar to what happens in bosonic supergravity between the Ricci scalar, $\hat{R}$, and the $\frac{1}{12}\hat{H}^{2}$ term, under the same limit. In this work we present the explicit form of the gauge transformations and gauge curvatures after considering a NR limit. We compute the full finite Lagrangian in its gauge covariant form, using gauge covariant curvatures and gauge covariant derivatives. These results constitute the first computation of a finite Lagrangian for heterotic supergravity under these kinds of limits. As an interesting feature of the expansion implemented, the Green-Schwarz mechanism for the two-form can be trivialized. Terms equivalent to Chern-Simons contributions naturally arise from the previous property.
@@ 1 Introduction
In the last decade, there has been significant progress made in the study of the different formulations of non-relativistic string theory [1]-[3] and its supergravity limit [4]-[17]  [^1] See [18]-[19] for reviews, and also [20] for a complementary introduction to this topic.. The low-energy limit of bosonic string theory is given by the universal NS-NS supergravity, whose fundamental fields $\hat{g}$, $\hat{B}$, $\hat{\phi}$ admit a non-relativistic expansion of the form [13],
@@
where $\mu,\nu=1,\dots,10$ are space-time indices and we have split the flat index $\hat{a}=(a,a^{\prime})$ where a,b,=0,1a,b,...=0,1 are the longitudinal directions and $a^{\prime},b^{\prime},...=2,\dots,9$. The quantities in the previous expansion obey the following Newton-Cartan relations,
@@
While, a priori, this expansion produces divergent tensorial quantities upon taking the NR limit $c\rightarrow\infty$ (i.e. Riemann tensor and its traces, H strength, etc.), the NS-NS supergravity Lagrangian,
@@
remains finite due to a curious cancellation between the divergent part of the Ricci scalar and the divergent part of the $\hat{H}$ squared term [13].
In the case of the bosonic sector of heterotic supergravity, we expect a similar cancellation when including the gauge field $\hat{A}_{\mu}^{i}$, $i=1,\dots,N$ and $N=496$ which is the dimension of the heterotic gauge group. Given a particular expansion for the gauge field, the inclusion of Chern-Simons terms in the curvature of the $\hat{B}-$field should produce divergences that must be compensated by the Yang-Mills Lagrangian. A first expansion for the $A$-field was constructed in [21], where the authors showed the cancellation of divergences due to the new gauge contributions. Another possible expansion was given in [22], using a manifest T-duality scenario, which also provides a finite supergravity guaranteed by T-duality invariance. Despite these recent efforts in this direction, there is still no explicit computation of the finite form of the heterotic Lagrangian under the NR limit, which is one of the goals of this work.
[…]
$\hat{\cal H}_{MN}$ $=$ $\hat{\cal H}^{(0)}_{MN}+\frac{1}{c^{2}}\hat{\cal H}^{(-2)}_{MN}\,$ (1.9)
$e^{-2\hat{d}}$ $=$ $e^{-2\hat{\Phi}}\sqrt{-\hat{g}}=e^{-2\phi}\sqrt{f(\tau,h)}=e^{-2d}\,$, (1.10)
where $f(\tau,h)=-\frac{\hat{g}}{c^{4}}$ and the different contributions to the generalized metric are given by
${\cal H}^{(0)}_{MN}$ $=\left(\begin{matrix}h^{\mu\nu}&\epsilon_{ab}\tau_{\nu}{}^{b}\tau^{\mu a}-b_{\rho\nu}h^{\rho\mu}\\ \epsilon_{ab}\tau_{\mu}{}^{b}\tau^{\nu a}-b_{\rho\mu}h^{\rho\nu}&h_{\mu\nu}+b_{\rho\mu}h^{\rho\sigma}b_{\sigma\nu}-2\epsilon_{cd}\tau_{(\mu|}{}^{d}b_{\sigma|\nu)}\tau^{\sigma c}\,\end{matrix}\right)$ (1.11)
${\cal H}^{(-2)}_{MN}$ $=\left(\begin{matrix}\tau^{\mu a}\tau^{\nu}{}_{a}&-b_{\rho\nu}\tau^{\rho c}\tau^{\mu}{}_{c}\\ -b_{\rho\mu}\tau^{\rho c}\tau^{\nu}{}_{c}&b_{\rho\mu}\tau^{\rho c}b_{\sigma\nu}\tau^{\sigma}{}_{c}\end{matrix}\right)\,.$ (1.12)
[…]
The extension from bosonic DFT to heterotic DFT can be easily done by extending the duality group from $O(D,D)$ to $O(D,D+N)$, where $N$ is the dimension of the heterotic gauge group. The components of the generalized metric in this case contain a metric $\hat{g}_{\mu\nu}$, a $\hat{B}-$field and a gauge field $\hat{A}_{\mu}{}^{i}$, which transforms as a connection under the gauge group. The $\hat{B}-$field inherits a non-covariant gauge transformation (Green-Schwarz mechanism), i.e.,
@@
with $\hat{\lambda}$ an arbitrary gauge parameter. In order to preserve the finiteness of the generalized metric when including the gauge field one possibility is to expand the field $\hat{A}_{\mu}{}^{i}$ following the proposal [22]
@@
where $v^{\pm}=\frac{1}{\sqrt{2}}(v^{0}\pm v^{1})$. The expansion (1.15) was constructed in such a way that the generalized metric of heterotic DFT, which is an element of $O(D,D+N)$, remains convergent after taking the NR limit in the double geometry. Demanding this condition from the very beginning allows one to have a finite action and equations of motion, by construction, upon considering the NR limit in the double geometry, since both the Lagrangian and the equations of motion depend on the generalized metric and dilaton (metric formulation of DFT). The construction of non-relativistic heterotic supergravity given in this work (which continues the ideas formulated in [22]) is an alternative formulation to the one presented in [21], and both constructions should be related by field redefinitions in order to have a unique supersymmetric extension (see [15] for the minimal supersymmetric extension).
@@
remains finite. We will also present the explicit form of the symmetry transformations and the full heterotic supergravity Lagrangian in its gauge covariant form upon considering the NR limit.
@@ 1.1 Organization of the paper
This paper is organized as follows: in the next section we review the cancellation of divergences in bosonic supergravity and we show the cancellation of divergences in heterotic supergravity. In section 3 we study the gauge transformation rules of the fundamental fields of the theory upon considering the NR limit and we construct a suitable gauge covariant derivative and gauge curvatures for the gauge fields. In section 4 we compute the finite Lagrangian of the theory in its gauge covariant form. Finally, in section 5, we present a discussion and an outlook.
@@ 2 Cancellation of divergences
When one expands the action (1.7), both the $\hat{R}$ and $\frac{1}{12}\hat{H}^{2}$ terms produce $c^{2}$ contributions. The same happens when we analyze (4.7), but this time the gauge compensation of $\frac{1}{12}\bar{H}^{2}$ is given also by the $\frac{1}{4}\hat{F}^{2}$ term. We will analyze both scenarios separately.
@@ 2.1 Ungauged case (review)
Let us review the construction given in [13], with a focus on the cancellation of divergences in the bosonic supergravity Lagrangian. To do so, we define
@@
[…]
So far we have just considered the expansion of the metric tensor (1.1) and its inverse (1.2). Therefore, for backgrounds where $h^{\mu\nu}$ is constant, the divergence in the Ricci scalar becomes trivial and it is possible to define a non-relativistic supergravity without including a non-relativistic expansion in the $\hat{B}-$field.
For a more general scenario we impose (1.3) and then the divergent part of $\frac{1}{12}\hat{H}^{2}$ is given by
[…]
Using the relation $\epsilon_{ab}\epsilon_{cd}\eta^{ac}=\eta_{bd}$ one can cancel the divergences coming from the Ricci scalar term.
@@ 2.2 Gauged case
upon considering the non-relativistic expansions (1.15) and (1.2), contains a divergence of the form (see appendix A for intermediate steps)
@@
These terms indicate that the $A$-expansion in (1.15) does not lead to a finite Yang-Mills theory. However, the inclusion of Chern-Simons terms in the heterotic Lagrangian (4.7) compensates the divergence (2.5). Indeed, one can compute (we have used CADABRA [31] for performing this test, but the main details of the cancellation are given in the Appendix A),
@@
At this point, we have proved that the heterotic supergravity Lagrangian is finite, but trying to write it in a non-covariant form leads to a huge number of terms. For this reason, our strategy will be to write it in a gauge covariant form, making use of the gauge invariance of the action.
In the next section, we will study the form of the gauge transformation rules and we will construct a suitable gauge covariant derivative. After that, we will construct new curvatures for all the fundamental fields transforming in a non-covariant way under gauge transformations.
@@ 3 Transformations and curvatures
In this section we will focus on the non-Abelian gauge transformations, in order to be able to write a gauge covariant action using tensors under this symmetry. We will start by describing the Levi-Civita connection, then we will construct the gauge covariant derivative and finally we will include gauge transformations with a gauging in the covariant derivative. Remarkably, we will describe the Green-Schwarz mechanism for the b-field and its trivialization, considering field redefinitions.
3.1 Levi-Civita connection
The Levi-Civita connection can be written as
$\Gamma_{\mu\nu}^{\rho}=c^{2}\Gamma_{\mu\nu}^{(2)\rho}+\Gamma_{\mu\nu}^{(0)\rho}+\frac{1}{c^{2}}\Gamma_{\mu\nu}^{(-2)\rho}\,$, (3.1)
where the explicit form of the previous terms is given by
$\Gamma_{\mu\nu}^{(2)\rho}$ $=$ $h^{\rho\sigma}(\partial_{\mu}\tau_{\sigma}{}^{a}-\partial_{\sigma}\tau_{(\mu}{}^{a})\tau_{\nu)}{}^{b}\eta_{ab}\,$, (3.2)
[…]
$\Gamma_{\mu\nu}^{(-2)\rho}$ $=$ $-\partial_{(\mu}\tau^{\sigma a}h_{\nu)\sigma}\tau^{\rho}{}_{a}-\tau^{\rho a}\tau^{\sigma}{}_{a}\partial_{\sigma}h_{\mu\nu}\,.$ (3.4)
Inspecting the transformation rules under infinitesimal diffeomorphisms given by the Lie derivative $L_{\xi}$, i.e.,
$\delta_{\xi}h_{\mu\nu}$ $=$ $L_{\xi}h_{\mu\nu}=\xi^{\rho}\partial_{\rho}h_{\mu\nu}+2\partial_{(\mu}\xi^{\rho}h_{\nu)\rho}\,$, (3.5)
$\delta_{\xi}h^{\mu\nu}$ $=$ $L_{\xi}h^{\mu\nu}=\xi^{\rho}\partial_{\rho}h^{\mu\nu}-2\partial_{\rho}\xi^{(\mu}h^{\nu)\rho}\,$, (3.6)
$\delta_{\xi}\tau_{\mu}{}^{a}$ $=$ $L_{\xi}\tau_{\mu}{}^{a}=\xi^{\rho}\partial_{\rho}\tau_{\mu}{}^{a}+\partial_{\mu}\xi^{\rho}\tau_{\rho}{}^{a}\,$, (3.7)
$\delta_{\xi}\tau^{\mu}{}_{a}$ $=$ $L_{\xi}\tau^{\mu}{}_{a}=\xi^{\rho}\partial_{\rho}\tau^{\mu}{}_{a}-\partial_{\rho}\xi^{\mu}\tau^{\rho}{}_{a}\,$, (3.8)
where $\xi^{\mu}$ is an arbitrary parameter. It is straightforward to check that $\Gamma_{\mu\nu}^{(2)\rho}$ and $\Gamma_{\mu\nu}^{(-2)\rho}$ transform covariantly under infinitesimal diffeomorphisms and they can be interpret as covariant shifts of the non-covariant part of the non-relativistic connection, given by $\Gamma_{\mu\nu}^{(0)\rho}$.
@@ 3.2 Covariant derivative
$=$ $\partial_{\mu}\tau_{\nu}{}^{a}+T_{\mu\nu}^{a}\,$,
where $T_{\mu\nu}{}^{a}=T_{\nu\mu}{}^{a}$ is given by the following quantity,
$T_{\mu\nu}{}^{a}$ $=$ $+\partial_{(\mu}{\tau_{\nu)}{}^{a}}+\eta_{bc}\eta^{ad}\nabla_{(\mu|}{\tau_{\rho}{}^{b}}\tau_{|\nu)}{}^{c}\tau^{\rho}{}_{d}-\eta_{bc}\eta^{ad}\nabla_{\rho}{\tau_{(\mu|}{}^{b}}\tau_{|\nu)}{}^{c}\tau^{\rho}{}_{d}$ (3.10)
$-\frac{1}{c^{2}}\partial_{(\mu}{\tau_{\nu)}{}^{a}}-\frac{1}{c^{2}}\eta_{bc}\eta^{ad}\nabla_{(\mu}{\tau_{\rho}{}^{b}}\tau_{\nu)}{}^{c}\tau^{\rho}{}_{d}+\frac{1}{c^{2}}\eta_{bc}\eta^{ad}\nabla_{\rho}{\tau_{(\mu}{}^{b}}\tau_{\nu)}{}^{c}\tau^{\rho}{}_{d}\,$.
An interesting observation of $\nabla_{\mu}\tau_{\nu}{}^{a}$ is that this quantity does not diverge, but contains $\frac{1}{c^{2}}$ terms that, in general, one needs to track in order to correctly write a finite Lagrangian. Moreover, the antisymmetric part of $\partial_{\mu}\tau_{\nu}^{a}$ satisfies
$\nabla_{[\mu}\tau_{\nu]}^{a}$ $=$ $\partial_{[\mu}\tau_{\nu]}{}^{a}\,$. (3.11)
Therefore, the antisymmetric part of $\partial_{\mu}\tau_{\nu}^{a}$ is covariant under diffeomorphims, giving rise to the well-known intrinsic torsion of the theory [13]. In order to write the two-derivative ungauged action in a covariant way we will just need $R^{(0)}$ given in (LABEL:Scalar) (which we prove in the appendix that it is a covariant quantity), and the relation (3.11) to replace ordinary derivatives by covariant derivatives in some antisymmetric expressions in the terms coming from the $\bar{H}^{2}$ contribution.